OpenStudy (anonymous):

please help: Use the substitution method to solve the linear system. -x+3y=4 x+6y=14 different problem 3x+2y=8 x+4y=-4

OpenStudy (anonymous):

Choose one equation from each pair and solve for y. Then substitute that y into the second equation. IE 3x+y=4 x+2y=7 Solving for y the first equation becomes y=-3x+4. Plug that in for y in the second equation: x+2(-3x+4)=7 Simplify for x. -5x=1 x=-1/5 choose one of original equations and plug in x and solve for y. 3(-1/5)+y=4 y=23/5

OpenStudy (anonymous):

-x+3y=4 x+6y=14 Take an equation: x + 6y = 14 Then make it equal x. x + 6y = 14 - 6y -6y x = - 6y + 14 Then plug x into the first equation that was not used: -(- 6y + 14) + 3y = 4 6y - 14 + 3y = 4 9y - 14 = 4 +14 +14 9y = 18 /9 /9 y = 2 Then plug y into either equation -x+3y=4 -x + 3(2) = 4 -x + 6 = 4 -6 -6 -x = -2 /-1 /-1 x = 2 The solution for the first problem is x = 2 and y = 2